### v4

# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

Abstract
Various industries such as finance, meteorology, and energy produce vast amounts of heterogeneous data every day. There is a natural demand for humans to manage, process, and display data efficiently. However, it necessitates labor-intensive efforts and a high level of expertise for these data-related tasks. Considering large language models (LLMs) showcase promising capabilities in semantic understanding and reasoning, we advocate that the deployment of LLMs could autonomously manage and process massive amounts of data while interacting and displaying in a human-friendly manner. Based on this, we propose Data-Copilot, an LLM-based system that connects numerous data sources on one end and caters to diverse human demands on the other end. Acting as an experienced expert, Data-Copilot autonomously transforms raw data into multi-form output that best matches the user’s intent. Specifically, it first designs multiple universal interfaces to satisfy diverse data-related requests, like querying, analysis, prediction, and visualization. In real-time response, it automatically deploys a concise workflow by invoking corresponding interfaces. The whole processes are fully controlled by Data-Copilot, without human assistance. We release Data-Copilot-1.0 using massive Chinese financial data, e.g., stocks, funds and news. Experiments indicate it achieves reliable performance with lower token consumption, showing promising application prospects.
Figure 1: Data-Copilot is an LLM-based system for data-related tasks, bridging billions of data and diverse user requests. It designs versatile interface tools for data management, invocation, processing, and visualization. Upon receiving a complex request, Data-Copilot autonomously invokes self-design interfaces and constructs a workflow to fulfill human intent. Lastly, raw data is converted into graphics, tables, and text forms tailored to human tastes.
1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. This wide-ranging, multiform data encapsulates critical insights that could be leveraged for a host of applications, from predicting financial trends to monitoring energy consumption.
Recently, the advancement of large language models (LLMs) (Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023), particularly the ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023a), has paved the way for advanced AI systems. Leveraging chain-of-thought (Wei et al., 2022c; Kojima et al., 2022; Gao et al., 2022b; Wang et al., 2023c), reinforcement learning from human feedback (Ouyang et al., 2022; Wang et al., 2022b), and instruction-following (Wei et al., 2022a; Iyer et al., 2022; Chung et al., 2022), LLMs demonstrate remarkable abilities in dialogue, reasoning and planning. However, in the face of the sheer magnitude and complexity of data, LLMs are confronted with the colossal challenge of managing, processing, and displaying data.
To achieve this, several challenges must be addressed: (1) From a data perspective: employing LLMs for directly reading and processing massive data is not only impractical but also poses the potential risks of data leakage. (2) From the model perspective: due to limitations of the input context, LLMs are unable to process large-scale data at once and are also not adept at handling numerical computations. (3) From the task perspective: many data-related tasks are intricate and require a combination of many operations, like retrieval, computations, and table manipulations, and present their results in multiple formats, e.g., images, tables, and text, which exceeds the capabilities of LLMs. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have embarked on valuable explorations. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023) and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL. However, our focus is more on how to better process and present data based on human preference, rather than merely retrieving it from databases. Additionally, while SQL is convenient, it can not directly satisfy common data analysis needs such as prediction and visualization.
In light of this, we contemplate how to integrate LLMs into every facet of data-related tasks, not just as a Text2SQL solver. More specifically, we explore how LLMs can efficiently and autonomously process vast amounts of data to meet diverse human needs, without the necessity of direct data access. To pursue the answer, we trace back to the origins of data science in the 1970s. The pioneer of data science, Peter Naur (Turing Award winner in 2005), defined data science as follows:
Data science is the science of dealing with data and processing large amounts of data. Humans as sources and receivers of data. The data must be chosen with due regard to the transformation to be achieved and the data processing tools available. —Naur, 1974
His statement elucidates how humans accomplish data-related tasks: humans create various tools that are easy to use and then transform the data into a human-friendly format. Similarly, LLMs also should play a more intelligent role like humans: designing tools for data management, and invoking tools for processing and presenting data.
Additionally, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). These LLM-based agents showcase the potential of LLMs in meeting human needs through code as an intermediary.
Inspired by them, we advocate utilizing the coding capabilities of LLMs to create diverse and powerful tools, thereby facilitating the management and utilization of vast amounts of data. Based on this, we propose Data-Copilot, a reliable data agent that manages vast data and dynamically fulfills a wide spectrum of human needs. As shown in Figure 1, Data-Copilot autonomously designs versatile interface tools for data management, and dispatch these interface tools step by step for diverse human requests, e.g., data analyzing, forecasting, and visualization, forming a data-to-human workflow.
As an LLM-based agent, Data-Copilot is versatile. It seamlessly connects an array of data sources from various domains on one end, while adeptly catering to diverse user demands on the other. Furthermore, Data-Copilot is scalable. It can continuously extend its capability boundaries by developing new interface tools for emerging new data and requests. Our system comprises two processes:
  • Interface Design: Data-Copilot adopts an iterative self-request process to fully explore the data, mimicking a broader range of human need scenarios. As shown in Figure 1, starting from a few seed requests, it self-generates a large number of requests and then abstracts them into generalized interface tools for efficient utilization. These interfaces empower Data-Copilot to efficiently perform data querying, processing, forecasting, and visualization.
  • Interface Dispatch: When a user request is received, Data-Copilot first parses the user intention and then plans an interface invoking workflow for the user request. This step-by-step invocation can efficiently fulfill user requirements while minimizing token overhead.
Figure 2: Overview of Data-Copilot. Interface Design: We devise a self-request process allowing LLM to generate sufficient requests from a few seed requests autonomously. Then LLM iteratively designs and optimizes interfaces based on generated requests. These interfaces are described in natural language, making them easily scalable and transferable across different platforms. Interface Dispatch: Upon receiving user requests, LLM plans and invokes interface tools based on self-design interface descriptions and in-context demonstrations. This allows for the deployment of a logical workflow that fulfills user demands and presents the results to the user in multi-form.
Incorporating two phases, Data-Copilot achieves autonomous data management and utilization without the need for direct data access. Our contributions can be summarized as follows:
  1. 1.
    To efficiently handle data-related tasks, we design a universal system, Data-Copilot, that connects massive data sources to diverse user tastes. It leverages LLM for more efficient data management and utilization, reducing tedious labor.
  2. 2.
    Data-Copilot can autonomously manage, process, analyze, and visualize data. For a human request, it transforms raw data into informative results that best match the user’s intent.
  3. 3.
    As a reliable assistant, Data-Copilot abstracts a multitude of common needs into various practical interface tools. When faced with real-time requests, it invokes these related interfaces sequentially or in parallel for problem-solving.
  4. 4.
    We evaluate Data-Copilot on massive Chinese financial data, e.g., stocks, funds, macroeconomics, and news. The excellent success rate and lower computational cost showcase its potential and application prospects.
2 Data-Copilot
Data-Copilot is an LLM-based agent that accomplishes automated management, invocation, and processing of a large number of data by self-design sophisticated interface tools. Besides, Data-Copilot adeptly comprehends user intentions, and invokes well-designed tool interfaces for request solving, significantly reducing the need for tedious labor.
As shown in  Figure 2, the LLM plays the role of a designer (Section 2.1) to independently design various tool interfaces to for data management and utilization. Besides, LLM also acts as a dispatcher (Section 2.2), invoking generated interfaces and deploying the workflow for problem-solving. Through the integration of two stages, Data-Copilot manages to efficiently handle a large volume of data-related tasks and caters to various user needs.
2.1 Interface Design
Just as Peter Naur pointed out, designing a set of versatile interface tools that are easy to manage and use is critical. Therefore, Data-Copilot designs a plethora of interfaces as tools, where each interface is composed of natural language (functional description) and code (implementation), responsible for data acquisition, processing, and visualization.
As shown in Figure 2, beginning with some seed requests, LLM autonomously mimics and generates a large number of requests. Then LLM designs common interfaces for these requests, representing a set of universal needs. By gradually refining and expanding their functionalities, these interfaces progressively cover a more diverse array of needs. Ultimately, the LLM translates these interfaces into executable and callable code modules.
Explore Data by Self-Request
The design of the interface depends on two aspects: what kind of data is available and also what kind of demand the user proposes. Inspired by (Wang et al., 2022a; Dibia, 2023), Data-Copilot first autonomously explores the data to mine more requests. Then it proceeds to design the interfaces for these requests.
Specifically, we generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. This process does not require too much labor, as data providers usually offer a description of the data and access methods.
Then we adopt a self-request process to explore data, i.e. the parsing files and a few seed requests are fed into the LLM and then LLM is instructed to generate more diverse requests by mimicking. This process aims to explore an extensive array of data sources while covering diverse user requests. In Figure 2, LLM generates four new requests based on the two seed requests and data parsing files. The detailed prompts are shown in Section F.1.
Interface Definition
In this step, Data-Copilot creates various interface tools to fulfill the previously generated requests. Specifically, we feed all data parsing files and all interfaces stored in the interface library (empty at first iteration) into LLM as a prompt. LLM needs to resolve each request one by one. When addressing each request, LLM can use existing interfaces from the interface library or redefine a new interface to fulfill the task.
Concretely, if existing interfaces are insufficient to address the current request, LLM can create a new interface: Interface1={Name:"", Function description:"", Input and Output:""}... After that, LLM proposes a feasible solution using these self-design or existing interfaces: Solution="These three interface functions will meet your needs. You should first use the interface: {getBankData} to acquire data, then...". The prompts and cases are shown in Sections F.1 and G.1.
Importantly, Data-Copilot only uses natural language to define the interface’s functions and their arguments, without considering the specific implementation, which is grammar-free. This process is similar to software architecture design by a human programmer. It allows Data-Copilot to focus on designing the layout and relation of the interfaces with different functionalities, rather than complying with programming grammar.
Interface Merging
To make the interface more universal, Data-Copilot considers whether each newly designed interface can be merged with the existing ones in the library. Specifically, when a new interface is designed, Data-Copilot also checks whether this interface is similar to the previous ones, in terms of functionality, arguments, etc. Two similar interfaces are merged to create a new and generalized interface. As shown in Figure 2, two interfaces {Interface: Query-GDP} and {Interface: Query-CPI} are merged into one {Interface: Query-GDP-CPI}. This merging process is similar to software developers wrapping similar modules into a new one.
After this process, a large number of similar interfaces are merged. Each interface within the library possesses distinct functionalities, which are adeptly suited for real-time response applications.
Interface Implementation
Data-Copilot iteratively processes each request through interface definition and interface merging. Eventually, when all the requests can be satisfied by these self-design interfaces, Data-Copilot transforms these interface tools into executable and callable code modules.
The whole interface design process is offline. As shown in Figure G3, Data-Copilot automatically produces five types of interfaces in the interface library: data acquisition, processing, prediction, visualization, and DataFrame manipulation. It transforms repetitive and tedious labor into an automated process, and also can effortlessly add new interfaces for additional requests or new data sources. Besides, Data-Copilot easily switches to other programming platforms and databases by simply re-generating the implementation code for the interface, demonstrating excellent scalability.
2.2 Interface Dispatch
In the first stage, Data-Copilot designs a variety of generic interface tools for data acquisition, processing, and visualization. These interfaces are versatile and easy to use, improving the efficiency of real-world response. These interfaces are invoked in this phase and form an efficient workflow for real-time requests (Figure 2). A detailed prompt and cases are shown in Sections F.2 and G.2
Intent Analysis
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, if the question is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2017 to May 2023 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this external information into LLM along with the original request to generate the parsed result.
Planing Workflow
Once the user’s intent is accurately understood, Data-Copilot plans a reasonable workflow to process the user’s request. We feed the user request and related interface descriptions into LLMs for the automatic deployment of workflows. Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot, based on the user’s request and the interface.
To achieve this, Data-Copilot generates a response in JSON format representing each step of the scheduling: ###Interface Call: {step1={"arg":"", "function":"", "output":"" ,"description":""}, step2={..}, ..}. Please refer to Section F.2 for a detailed prompt. Data-Copilot meticulously orchestrates the scheduling of the interface in either a sequential or parallel manner.
For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot first plans five steps in its workflow. In the first two steps, it dispatches {Data Acquisition}, {Calcuate Return} in parallel for three indices. The last three steps successively invoke the {Visualization} to sequentially plot three indices on the same canvas. In the second case, Data-Copilot deploys a loop workflow to sequentially calculate the indicators of 50 constituent stocks by calling the {Loop and Rank}.
Multi-form Output
Upon the deployment and execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. This systematic summary not only provides clear and concise results but also sheds light on the steps taken to achieve those results, thereby enhancing transparency and understanding of user requests.
As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent based on local time. Then, it deploys a three-step workflow: 1) Invoking {get-GDP-data} interface to acquire historical GDP data. 2) Invoking {predict-next-value} interface for forecasting. 3) Visualizing the output.
3 Qualitative Evaluation
3.1 Environment and Data Sources
We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies. The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance. In the first phase, we use the gpt-4-0613222https://platform.openai.com/ for interface design and gpt-3.5-turbo-0613 for interface dispatch.
3.2 Seeds and Datasets
Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each. These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate. 2) involves different levels of difficulty. Then we filter out highly similar requests by computing the sentence embedding of each request. Lastly, we retain 173 high-quality requests, which are used as the seeds for Interface Design phase of the Data-Copilot.
Self-Request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request strategy to expand the request set. Specifically, we feed 173 manually proposed seed requests along with all data descriptions and schema into GPT-4. We instruct GPT-4 to generate massive requests, covering diverse data. In this phase, we generate a total of 6480 requests. Then we compute the semantic similarity between all requests and filter out those with high similarity. Lastly, we retain 3547 requests. We randomly select 3000 requests from the total set for subsequent interface design. Meanwhile, the remaining 547 requests are treated as a test set. We manually annotate the final answers, task types, and workflow types for each testing instance. The statistics of three dataset are shown in Tables 1 and C1.
3.3 Qualitative Results
The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request. The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces. For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
Data-Copilot designs many versatile interface tools through a self-request. As shown in Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, dataframe manipulation, and visualization. These diverse interfaces endow the LLM with the capability to address complex requests. For instance, get-stock-prices-data interface integrates the functions of acquiring data in daily, weekly, and monthly frequencies, controlled by argument: freq, which is very similar to the process of programmer development.
Data-Copilot deploys workflows with different structures. Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests. As shown in Figures G4 and G5, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface. It iteratively calculates the financial indicators for each stock. Please refer to Appendix E for more visualization cases.
4 Quantitative Evaluation
We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
Table 1: Left: Statistics on human-proposed requests and Self-Request dataset. Right: Detail Statistics on different request types and workflow structures.
Source #Cases Form Usage
Human-proposed 173 Query Seed
Self-Request 3000 Query Interface Design
Self-Request + Human-annotated 547 Query, Ans Testing
Request Types Workflow Types
Sequential Parallel Complex Overall
Stock 79 106 87 272
Fund 55 36 22 113
Corp. 97 30 4 131
Other 6 12 13 31
Total 237 184 126 547
4.1 Experiments Settings
Baselines 1. Global Planning (GP) (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests. 2. ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step. 3. Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 4. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. All methods are evaluated on the 547 testing instances (Section 3.2) three times. A detailed prompts of baselines is shown in Section F.3.
Metrics We design three metrics for solution code and answer: Exec@1 Pass@1 and Efficiency. Exec@1 evaluates whether the code can be executed by a compiler without errors. Pass@1 measures the correctness of the final answer. Considering the diverse output formats, e.g., images, text, numbers, and tables, we employ GPT-4 to automatically assess the answer. If the output is an image, we use gpt-4-vision-preview to compare the generated image to the label. Lastly, we calculate the total number of tokens consumed by the input prompt and output response, which denotes the solving efficiency of each method.
Table 2: All methods use gpt-3.5 as based LLM.
Methods Exec@1(%) Pass@1(%) #Token
GP 50.8 28.5 823.6
ReAct 63.4 44.1 1515.9
Reflexion 75.7 59.0 2463.5
Multi-Agent 73.6 57.4 2835.2
Data-Copilot 79.9 70.2 561.2
Table 3: All samples are divided into three subsets based on its complexity. We compute Pass@1.
Methods Sequential Parallel Complex Overall
GP 42.4% 29.0% 1.6% 28.5%
ReAct 56.3% 48.6% 14.6% 44.1%
Reflexion 67.8% 55.1% 48.1% 59.0%
Multi-agent 63.2% 64.7% 35.5% 57.4%
Data-Copilot 71.8% 70.0% 67.1% 70.2%
4.2 Comparison Results
As shown in Table 3, Data-Copilot surpasses all baselines in success rate (Pass@1), Exec@1, and Efficiency, achieving 11.2% improvements.
Data-Copilot significantly reduces the risk of failure. Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%. In most cases, our system merely needs to invoke pre-designed interface tools and generate corresponding arguments to address user requests, significantly reducing the risk of errors. In contrast, baseline strategies must construct all code from scratch, which makes it easy to overlook certain details or make mistakes. Besides, the difference between Pass@1 and Exec@1 also indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
Data-Copilot reduces repetitive overhead for common requests. The consumed tokens represent computational overhead. Data-Copilot only consumes 561 tokens, which is just 30% of the consumption compared to baselines. Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments. The primary token consumption lies in the input prompt which contains related tool descriptions. In contrast, baseline strategies may consume a substantial amount of tokens since they need to output a complete solving code for each request. In the real world, numerous requests are similar or even repetitive. Data-Copilot can reduce repetitive overhead by abstracting common demands into interface tools. However, baseline methods are compelled to repetitively generate code for each request.
4.3 Analysis
Data-Copilot exhibits a pronounced advantage on more challenging tasks. To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex. As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing. In contrast, our methods effectively handle these tasks by invoking versatile interfaces and deploying workflow. These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
We can deploy Data-Copilot locally using small LLMs with commendable performance. We analyze the performance of Data-Copilot with different LLMs. As shown in Figure 4, we test GPT family APIs and various open-source models. We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results. For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo). This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
4.4 Human Evaluation
To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness. Specifically, we collect the solutions of Data-Copilot and Reflexion. Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix C). We present results in Figure 4. It demonstrates that our solution significantly outperforms the competitor in terms of rationality. Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%. Besides, as for conciseness, Data-Copilot lags slightly behind its competitor, with 27.8% wins against 30% losses. This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are quite universal, accommodating a wide array of functions, which makes our system less concise. The results of human evaluation are consistent with the previous assessment. Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
Figure 3: Results of Data-Copilot with various LLMs.
LLM Pass@1 (%) #Token
gpt-4 75.6 561.2
gpt-3.5-turbo 70.2 559.3
Llama-2-70B-chat 49.3 623.6
codellama-13B 50.8 507.4
text-davinci-003 47.5 495.9
Llama-2-13B-chat 36.9 443.8
Vicuna-13B-v1.5 33.2 502.8
Figure 4: Win rate for Data-Copilot Vs. Reflexion on human preferences.
5 Conclusion
We propose a universal framework, Data-Copilot, to address a wide array of data-related tasks. It acts as a bridge between numerous heterogeneous data and humans, effectively managing, processing, and presenting data according to human tastes, reducing the dependence on tedious labor and expert knowledge. Acting like an experienced expert, Data-Copilot autonomously designs universal interface tools for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
6 Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62376245), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the project of the Donghai Laboratory (Grant no. DH-2022ZY0013), and MOE Engineering Research Center of Digital Library.
Appendix
Appendix A Discussion
A.1 What is Data-Copilot?
  • Data-Copilot is an efficient LLM-based data management agent, that processes massive data and rapidly responds to real-time human requests.
  • It contains interface design and invocation phases. Interface design involves a self-exploration phase for data sources to create many versatile interfaces.
  • For interface invocation, it parses human real-time requests and invokes the designed interfaces to retrieve and process relevant data and generates multi-form outputs, e.g., text, charts, and tables.
A.2 Why Data-Copilot is designed in two stages?
  • Reduces redundant generation: People often have a large number of similar or repetitive query needs, e.g., many people are interested in seeing which stock has the largest gain today. For these repetitive query requests, employing the LLM to write code from scratch for each user request would be highly inefficient and a significant waste of computational resources. To avoid inefficient and repetitive code generation for numerous similar requests, we abstract common demands into generic interfaces through data self-exploration in advance (Phase-1). For real-time requests (Phase-2), it only needs to invoke these interfaces, reducing redundant generation.
  • Performs better and consumes fewer tokens: Compared to a complete code, it only needs to invoke the relevant interfaces, which is simpler and less prone to errors. Our results show that it performs better and consumes fewer tokens.
  • Stronger interpretability: Compared to complex code, interface workflow is easier for humans to read and inspect.
Appendix B Experiments Details
During the interface design phase, we instruct LLM to generate Python code for each interface in the interface implementation. We filter out interfaces that cannot run properly at each iteration. For the interface Dispatch phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface. We set the temperature to 0.2 for more consistent results.
Appendix C Human Evaluation Details
We collect the solutions of Data-Copilot and Reflexion on 173 human-proposed requests. Subsequently, we involve experts to evaluate which of the two solving codes is more rational and concise. The criteria for expert are as follows:
Rationality:
1) Accurately identifies data of interest to the user.
2) Processes the data logically.
3) Output data in a format that meets human needs, with aesthetic and concise presentation.
Conciseness:
1) Only extracts and computes data relevant to the users interest.
2) Code is clear, concise, and well-structured.
3) The solution process is devoid of repetition and redundancy, ensuring a straightforward and intuitive approach.
Experts are required to evaluate which of the two solutions is better based on these principles, and then we calculate the win rate of Data-Copilot against the Reflexion.
Figure C1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; OpenAI, 2023b; Wei et al., 2022b; Wu et al., 2023c) have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism (Ouyang et al., 2022; Wang et al., 2022b; a).
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  (Schick et al., 2023; Gao et al., 2022a; Qin et al., 2023a; Hao et al., 2023; Qin et al., 2023b) integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality (Wu et al., 2023a; Surís et al., 2023; Shen et al., 2023; Liang et al., 2023b; Huang et al., 2023). In addition, there are many LLM-based agent applications (Xie et al., 2023), such as CAMEL (Li et al., 2023a), AutoGPT [^3] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^4] https://github.com/reworkd/AgentGPT, BabyAGI [^5] https://github.com/yoheinakajima/babyagi, BMTools [^6] https://github.com/OpenBMB/BMTools, LangChain [^7] https://github.com/hwchase17/langchain, Agentverse (Chen et al., 2023b), Autoagent (Chen et al., 2023a), MetaGPT (Hong et al., 2023), AutoGEN (Wu et al., 2023b), etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies (Cai et al., 2023; Qian et al., 2023) have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME (Joshi et al., 2023) investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT (Jiang et al., 2023) explore reasoning abilities of LLM over structured data. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Besides, many reseraches (Liu et al., 2023; Chang & Fosler-Lussier, 2023; Dong et al., 2023), like Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DAIL-SQL (Gao et al., 2023), DIN-SQL (Pourreza & Rafiei, 2023), DB-Copilot (Wang et al., 2023b), MAC-SQL (Wang et al., 2023a), ChatDB (Hu et al., 2023), SQL-PaLM (Sun et al., 2023), and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL and table rasoning. Chain-of-Table (Wang et al., 2024) proposes a step-by-step reasoning strategy based on the table.
Distinct from these approaches, our Data-Copilot system is different in the following aspects:
(1) Different Task Settings: Unlike Text2SQL methods, Data-Copilot focuses more on how to utilize data and present it according to user preferences, rather than retrieving data from databases. We assume that Data-Copilot has access to a vast amount of data, which is available via API.
(2) Richer Usage Scenarios: Data-Copilot needs to meet a diverse range of human needs, such as forecasting and plotting. Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
(3) Distinct Implementation Approach: Data-Copilot designs its tools for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.
Appendix E Visualization
We provide several cases in this section to visualize workflow deployed by Data-Copilot, which includes queries about diverse sources (stocks, company finance, funds, etc.) using different structures (parallel, serial, and loop Structure).
Different Structures
As shown in Figure G4,G5, Data-Copilot deploys different structural workflows based on user requirements. In Figure G4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel. In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas. These concise workflows can cope with such complex requests well, which suggests that the interface design and dispatch process of Data-Copilot are rational and effective.
Diverse Sources
Figure G6, G7, G8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc. Although the formats and access methods of these data types are quite different, our system efficiently manages and displays the data through its self-designed versatile interface, requiring minimal human intervention.
Appendix F Detailed Prompts
We provide detailed prompts for our Data-Copilot and baselines. Specifically, for the interface design, we provide detailed prompts in Section F.1 for four phases: Self-Request, Interface Definition, Interface Merging, and Interface Implementation. For the interface dispatch, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow. Additionally, we outline prompts for all baselines in Section F.3: Global Planning, ReAct, Reflection, and Multi-Agent Collaboration strategies.
F.1 Prompts For Interface Design
Explore Data by Self-Request Phase
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###Data files:{
GDP-Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: You can access the data by pro.cn-gdp(start-time, end-time, frequency,...), start-time means....,
Output Schema: The data return 9 columns, including quarter: quarter, gdp: cumulative GDP, gdp-yoy: quarterly Year-on-Year growth rate, pi,...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’, frequency=’quarter’),
First Row: {2019Q4 990865.1 6.10, ...},
Last Rows: {2018Q4 900309.5 6.60, ...}
stock_basic:...,
Interface Definition Phase:
###Instruction: You are an experienced program coder. Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces (similar to pseudocode) to resolve my request.
(1) You should define the name, function, inputs, and outputs of the interface. Please describe the functionality of the interface as accurately as possible and do not write any implementation code in the new interface, just design the interface as existing interfaces.
(2) Finally please explain how to resolve my request using your newly designed interfaces or existing interfaces in the neural language.
###Output Format:
Your newly designed interfaces:
Interface1={Interface Name: {name},
Function description: {This interface is to ...},
Input: {argument1: type, argument2: type, ...},
Output: {pd.DataFrame} }
Interface2=....
The solving process for request: {To fullfil this request, I design a interface ...}
###The user request: {input request}
###Data Files: {All data files}
###The existing interfaces: {all interfaces in library}
Interface Merging Phase
###Instruction: Please check that the interface you have designed can be merged with any existing interfaces in the library.
(1) You should merge interfaces with similar functionality and similar input and output formats into a new interface.
(2) You can use parameters to control the different inputs. Finally
(3) Please explain your reason for merging and output all interfaces in the library after merging.
(4) If you dont think a merge is necessary, then just add new interfaces into the existing interface library and output them all.
###New interfaces: {interfaces}
###Existing interfaces: {interface1, interface2, interface3, ..}
###Output Format:
The reasons for merging: {reason}
All interfaces after merging: {interface1, interface2, merged interface3, ...}
Interface Implementation Phase:
###Instruction: Please generate the complete code according to the definition of a given interface and data access methods. The generated code must be in strict accordance with the definition of the interface and the formatting requirements of the arguments.
###Interface: {Input Interface}
###Data Files: {All data files}
F.2 Prompts For Interface Dispatch
Intent Analysis Phase
### Analysis Prompt: Please parse the input request for time, place, object, and output format. You should rewrite the instruction according to todays date. The rewritten new instruction must be semantically consistent and contain a specific time and specific indicators.
### Output Format: Rewritten Request. (Time:%s, Location:%s, Object:%s, Format:%s).
### User Request: Today is {Timestamp}. The user request is {Input Request}.
Please output a Rewritten Request.
Task Selection
###Select Prompt: Please select the most suitable task according to the given Request and generate its task_instruction in the format of task={task_name: task_instruction}. There are four types of optional tasks. [fund_task]: used to extract and process tasks about all public funds. [stock_task]: for extracting and processing tasks about all stock prices, index information, company financials, etc., [economic_task]: for extracting and processing tasks about all Chinese macroeconomic and monetary policies, as well as querying companies and northbound funds, [visualization_task]: for drawing one or more K-line charts, trend charts, or outputting statistical results.
###Output Format: task1={%s: %s}, task2={%s: %s}
###User Request: {Rewritten Request}.
Please output a task plan for this request.
Planning Workflow
###Planning prompt: Please use the given interface (function) to complete the Instruction step by step. At each step you can only choose one or more interfaces from the following interface library without dependencies, and generate the corresponding arguments for the interface, the arguments format should be strictly in accordance with the interface description. The interface in the later steps can use results generated by previous interfaces.
###Output Format:
Please generate as json format for each step:step1={"arg1": [arg1,arg2...], "function1": "%s", "output1": "%s", "description1": "%s"}, step2={"arg1": [arg1,arg2..], "function1": "%s", "output1": "%s", "description1": "%s"}, ending with ###.
###User Request: {Task Instruction}.
Please output an interface invocation for this instruction.
F.3 Prompts For Baselines
Global Planning
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, you should write a complete Python code to fulfill the users request. Your code must completely fulfill all the users requirements without syntax errors!
###User Request: {User request}
###Data files: {All data files}
Please solve the request by Python Code.
Step-by-Step ReAct
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and generate your thoughts and actions for each step, and then finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please summarize the results of the code execution just now and think about whether this result accomplishes what you planned for this step.
### {Action: Yes, I observed that this function successfully fetched the data...}
Step-by-Step Reflexion
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and then generate your thoughts and actions for each step. After the execution of your current action, you need to reflect on the results until your current plan has been successfully completed. Then you think about the next step and then generate your next action, and finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please record the compilers return results just now and think about whether this result accomplishes what you planned for this step.
### {Action: No, I observe that the compiler returns an error...}
### Reflection Prompt: Please reflect on the error returned by the compiler and regenerate a new Python code to resolve the issue. If the compilation passes without any errors, reflect on whether the current result is what you planned to do.
### {Action: I revise my solution as follows: def get-data2()....}
Multi-agent collaboration
###Instruction For Manager: You are the manager of the project team and you need to lead your team to fulfill user requests. You have two experienced programmers under you (Programmer-A and -B) and you need to assign them the same or different tasks according to the user request, then organize the discussion, and finally solve the problem.
###Instruction For Agent1/Agent2: You are an experienced programmer. Your team has a colleague who is also a programmer and a manager. You need to write code according to the managers arrangement, discuss it with them, improve your program, and get a consensus conclusion.
###User Request: {User request}
###Data files: {All data files}
-------- Phase1 Discussion for Task Assignment -------
###Task Assignment Prompt For 3 Agents: Now lets start discussing how to solve user problems. The speaking order is manager, A and B. After a round of discussion, the manager decides on a preliminary solution and assigns specific tasks to the two programmers
###Manager Response: {To fulfill the request...}
###Programmer A Response: {Yes, I agree with the manager. We should first,....}
###Programmer B Response: {That’s a good solution. We also should...}
###Manager Response: {After discussion, Programmer A should..., B should ....}
-------- Phase2 Coding and Discussion-------
###Prompts For 3 Agents: Two programmers complete their respective tasks as assigned by their manager. Then the manager organizes a discussion on whether the two options and their running results are reasonable. Lastly, two programmers need to regenerate the code.
###Programmer A Response: {Solution1 Generated by A}
###Programmer B Response: {Solution1 Generated by B}
###Prompts For 3 Agents: Please evaluate two solutions. {Solution1, Result1, Solution2, Result2}
###Programmer A Evaluation: {I think ...}
###Programmer B Evaluation: {I agree with ...}
###Programmer A Regeneration: {Revised Solution1}
###Programmer B Regeneration: {Revised Solution2}
-------- Phase3 Discussion-------
-------- Phase4 Final Decision-------
###Prompts For manager: Please output a complete and correct solving code for user request based on the previous discussions of the group.
###Manager Decision: {Final Solution}
Appendix G Case Study
G.1 Cases for Interface design
---- The First Request: Interface Definition Phase -----
Input Request: Get the daily opening price data of Bank of Hangzhou from Jan 23, 2018 to Mar 13, 2019 and plot the trend.
Response: First, we need to design a function to get the opening price data from the banks database. Then, we need another function to process this data and prepare them for graphical representation. Finally, we need a function to plot the trend chart.
New Design Interfaces:
{Interface1=Interface Name:getBankData, Function:.., Input:.., Output:..}
{Interface2=Interface Name:prepareDataForPlotting, Function:.., Input:.., Output:..}
{Interface3=Interface Name:plotData, Function:.., Input:.., Output:..}
The solving process for request: These three new interface functions will meet your needs. In your program, you would first get the data using getBankData, then process the data using prepareDataForPlotting, and finally use plotData to plot the trend.
---- The Second Request: Interface Definition Phase-------
Input Request: Get the monthly closing price data of Ningde Times from January 23, 2018 to March 13, 2019, and plot the trend.
Response: To fulfill this requirement, we design a new interface function. In addition, I need to use two existing interfaces to fulfill your requirement.
New Designed Interfaces:
{Interface4=Interface Name: getStockData..}
Existing Interfaces
{Interface1=Interface Name:getBankData, ...}
{Interface2=Interface Name:prepareDataForPlotting, ...}
{Interface3=Interface Name:plotData, ...}
The solving process for request: You need to call these functions to meet your needs. First, use new interface {getStockData} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting} to process this data, and finally use {plotData} to plot the trend...
---- Interface Merge Phase -------
The reason for merging: Yes, these interface functions can be merged and reformulated to eliminate duplicate functions. We can merge getBankData and getStockData into a more generic function get-stock-data, which can fetch any financial data, including bank and stock data. In addition, we can keep the prepareDataForPlotting and plotData functions, as they can also be applied to different types of financial data]
{Interface1=Interface Name:get-stock-data, ...}
{Interface2=Interface Name:prepareDataForPlotting, ...}
{Interface3=Interface Name:plotData, ...}
---- Interface Implementation -------
Using get_stock_data as examples:
def get_stock_data(name:str=’’,start_d:str=’’,end_d:str=’’,freq:str)->pd.DataFrame:
Retrieves the daily/weekly/monthly price data for a given stock code during a specific time period.
Args:
-name (str): The name of the stock.
-start_d (str): The start date in the format YYYYMMDD’.
-end_d (str): The end date in YYYYMMDD’.
-freq (str): The frequency of the price, can be daily’, weekly’, or monthly’.
Returns:
-pd.DataFrame: A dataframe that contains the daily/weekly/monthly data. The output columns contain stock_code, trade_date, open, high, low, close, pre_close, change), pct_chg,vol,amount
stock_code = get_stock_code(name)
if freq == daily’:
G.2 Cases for Real-time Request
### Request: Today is 2019-03-13. The user request is to please help me plot the stock price trend of Guizhou Maotai from January 23, 2018 to today.
### Rewritten Request: Please help me plot the closing price of Guizhou Maotai from Jan 23, 2018 to Mar 13, 2019 (Time:20180123-20190313, Location: China, Object: Guizhou Maotai stock price, Format: Line).
###Plan: task1={"stock_task": "Get the time series data of Guizhou Maotais stock price from 20180123 to 20190313"}, task2={"visualization_task": "Plot the stock price of Guizhou Maotai from 20180123 to 20190313 on a line chart"}
###Function Call:
step1={
"arg1": ["Guizhou Maotai", "20180123", "20190313", "daily"],
"function1": "get_stock_prices_data",
"output1": "result1",
"description1": "Guizhou Maotai daily stock price (open, close, high, low, etc.) time series data"},
step2={
"arg1": ["result1","close"],
"function1": "calculate_stock_index",
"output1": "result2v,
"description1": "Guizhou Maotais closing price time series data"},
step3={
"arg": ["result2", null, "line", "Guizhou Maotai closing price chart from January 23, 2018 to March 13, 2019"], "function": "plot_stock_data",
"output": "result1",
"description": "Guizhou Maotai Stock Price Line Chart"
Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.

### v5

@@
 Abstract
-Various industries such as finance, meteorology, and energy produce vast amounts of heterogeneous data every day.
-There is a natural demand for humans to manage, process, and display data efficiently.
-However, it necessitates labor-intensive efforts and a high level of expertise for these data-related tasks.
-Considering large language models (LLMs) showcase promising capabilities in semantic understanding and reasoning, we advocate that the deployment of LLMs could autonomously manage and process massive amounts of data while interacting and displaying in a human-friendly manner.
-Based on this, we propose Data-Copilot, an LLM-based system that connects numerous data sources on one end and caters to diverse human demands on the other end.
-Acting as an experienced expert, Data-Copilot autonomously transforms raw data into multi-form output that best matches the user’s intent.
-Specifically, it first designs multiple universal interfaces to satisfy diverse data-related requests, like querying, analysis, prediction, and visualization.
-In real-time response, it automatically deploys a concise workflow by invoking corresponding interfaces.
-The whole processes are fully controlled by Data-Copilot, without human assistance.
-We release Data-Copilot-1.0 using massive Chinese financial data, e.g., stocks, funds and news.
-Experiments indicate it achieves reliable performance with lower token consumption, showing promising application prospects.
-Figure 1: Data-Copilot is an LLM-based system for data-related tasks, bridging billions of data and diverse user requests.
-It designs versatile interface tools for data management, invocation, processing, and visualization.
-Upon receiving a complex request, Data-Copilot autonomously invokes self-design interfaces and constructs a workflow to fulfill human intent.
-Lastly, raw data is converted into graphics, tables, and text forms tailored to human tastes.
+Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily.
+Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive.
+Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution.
+However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget.
+Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests.
+The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks.
+Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response.
+Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation.
+When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent.
+Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests.
+Additionally, interface workflows are more efficient and offer greater interpretability than code.
+We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects.
+Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
+Figure 1: We present two strategies for using LLMs to automate data analysis.
+Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints.
+Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks.
+It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests.
+Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
 1 Introduction
 In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others.
-This wide-ranging, multiform data encapsulates critical insights that could be leveraged for a host of applications, from predicting financial trends to monitoring energy consumption.
-Recently, the advancement of large language models (LLMs) (Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023), particularly the ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023a), has paved the way for advanced AI systems.
-Leveraging chain-of-thought (Wei et al., 2022c; Kojima et al., 2022; Gao et al., 2022b; Wang et al., 2023c), reinforcement learning from human feedback (Ouyang et al., 2022; Wang et al., 2022b), and instruction-following (Wei et al., 2022a; Iyer et al., 2022; Chung et al., 2022), LLMs demonstrate remarkable abilities in dialogue, reasoning and planning.
-However, in the face of the sheer magnitude and complexity of data, LLMs are confronted with the colossal challenge of managing, processing, and displaying data.
-To achieve this, several challenges must be addressed: (1) From a data perspective: employing LLMs for directly reading and processing massive data is not only impractical but also poses the potential risks of data leakage.
-(2) From the model perspective: due to limitations of the input context, LLMs are unable to process large-scale data at once and are also not adept at handling numerical computations.
-(3) From the task perspective: many data-related tasks are intricate and require a combination of many operations, like retrieval, computations, and table manipulations, and present their results in multiple formats, e.g., images, tables, and text, which exceeds the capabilities of LLMs.
+Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption.
+However, these data-related tasks often require specialized knowledge and tedious manual labor.
+Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks.
+Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
+An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text.
+However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text.
+Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources.
+(2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations.
+LLMs are not adept at performing these tasks.
+(3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data.
 These challenges significantly constrain the application of LLMs in data-related tasks.
-Recently, many works have embarked on valuable explorations.
+Recently, many works have explored alternative solutions.
 LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration.
-Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DIN-SQL (Pourreza & Rafiei, 2023) and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL.
-However, our focus is more on how to better process and present data based on human preference, rather than merely retrieving it from databases.
-Additionally, while SQL is convenient, it can not directly satisfy common data analysis needs such as prediction and visualization.
-In light of this, we contemplate how to integrate LLMs into every facet of data-related tasks, not just as a Text2SQL solver.
-More specifically, we explore how LLMs can efficiently and autonomously process vast amounts of data to meet diverse human needs, without the necessity of direct data access.
-To pursue the answer, we trace back to the origins of data science in the 1970s.
-The pioneer of data science, Peter Naur (Turing Award winner in 2005), defined data science as follows:
-Data science is the science of dealing with data and processing large amounts of data.
-Humans as sources and receivers of data.
-The data must be chosen with due regard to the transformation to be achieved and the data processing tools available.
-—Naur, 1974
-His statement elucidates how humans accomplish data-related tasks: humans create various tools that are easy to use and then transform the data into a human-friendly format.
-Similarly, LLMs also should play a more intelligent role like humans: designing tools for data management, and invoking tools for processing and presenting data.
-Additionally, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b).
-These LLM-based agents showcase the potential of LLMs in meeting human needs through code as an intermediary.
-Inspired by them, we advocate utilizing the coding capabilities of LLMs to create diverse and powerful tools, thereby facilitating the management and utilization of vast amounts of data.
-Based on this, we propose Data-Copilot, a reliable data agent that manages vast data and dynamically fulfills a wide spectrum of human needs.
-As shown in Figure 1, Data-Copilot autonomously designs versatile interface tools for data management, and dispatch these interface tools step by step for diverse human requests, e.g., data analyzing, forecasting, and visualization, forming a data-to-human workflow.
-As an LLM-based agent, Data-Copilot is versatile.
-It seamlessly connects an array of data sources from various domains on one end, while adeptly catering to diverse user demands on the other.
-Furthermore, Data-Copilot is scalable.
-It can continuously extend its capability boundaries by developing new interface tools for emerging new data and requests.
-Our system comprises two processes:
-  • Interface Design: Data-Copilot adopts an iterative self-request process to fully explore the data, mimicking a broader range of human need scenarios.
-As shown in Figure 1, starting from a few seed requests, it self-generates a large number of requests and then abstracts them into generalized interface tools for efficient utilization.
-These interfaces empower Data-Copilot to efficiently perform data querying, processing, forecasting, and visualization.
-  • Interface Dispatch: When a user request is received, Data-Copilot first parses the user intention and then plans an interface invoking workflow for the user request.
-This step-by-step invocation can efficiently fulfill user requirements while minimizing token overhead.
+Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL.
+Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b).
+These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
+Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent.
+Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM.
+However, creating a reliable code agent for various data analysis tasks is also not an easy feat.
+Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments.
+Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
+Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis.
+First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent.
+The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks.
+Besides, Data-Copilot also incorporates a data exploration phase.
+During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance.
+In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
+This design can bring three advantages.
+Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses.
+Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency.
+Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
+To achieve this, as shown in Figure 1, Data-Copilot contains three steps for data exploration:
+  1. 1.
+    Request Exploration: Data-Copilot explores potential requests and mines a broader range of human need scenarios from data.
+It involves a "self-instruct" process to generate massive simulated requests based on all data schemas (types and formats) and seed requests.
+  2. 2.
+    Interface Design: Our agent generates code for each request and then considers whether these requests have similar demands and whether general interfaces can be abstracted.
+  3. 3.
+    Interface Optimization: To improve the quality of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces by reflecting on feedback from the compiler.
+After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks.
+When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests.
+For different request scenarios, Data-Copilot can flexibly design different interface invocation procedures, such as serial workflows, i.e., step-by-step invocation, parallel workflows, and loop workflows.
+Our contributions are threefold:
 Figure 2: Overview of Data-Copilot.
-Interface Design: We devise a self-request process allowing LLM to generate sufficient requests from a few seed requests autonomously.
-Then LLM iteratively designs and optimizes interfaces based on generated requests.
-These interfaces are described in natural language, making them easily scalable and transferable across different platforms.
-Interface Dispatch: Upon receiving user requests, LLM plans and invokes interface tools based on self-design interface descriptions and in-context demonstrations.
-This allows for the deployment of a logical workflow that fulfills user demands and presents the results to the user in multi-form.
-Incorporating two phases, Data-Copilot achieves autonomous data management and utilization without the need for direct data access. Our contributions can be summarized as follows:
+Data Exploration: It devises a self-request process to generate numerous potential requests based on data schemas.
+Then it interleaves to design and optimize the interfaces to solve these requests.
+The whole exploration-design-optimization process is operated in advance, generating many general and error-free interface modules, such as data retrieval, computation, and visualization.
+Workflow Deployment: Upon receiving real-time requests, LLM invokes these interfaces, forming an interface workflow with different structures for request solving.
   1. 1.
-    To efficiently handle data-related tasks, we design a universal system, Data-Copilot, that connects massive data sources to diverse user tastes.
-It leverages LLM for more efficient data management and utilization, reducing tedious labor.
+    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks.
+It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
   2. 2.
-    Data-Copilot can autonomously manage, process, analyze, and visualize data. For a human request, it transforms raw data into informative results that best match the user’s intent.
+    We decouple code generation into two phases: data exploration and workflow deployment.
+In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data.
+These interfaces enhance the success rate and efficiency of real-time responses.
+In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
   3. 3.
-    As a reliable assistant, Data-Copilot abstracts a multitude of common needs into various practical interface tools.
-When faced with real-time requests, it invokes these related interfaces sequentially or in parallel for problem-solving.
-  4. 4.
-    We evaluate Data-Copilot on massive Chinese financial data, e.g., stocks, funds, macroeconomics, and news.
-The excellent success rate and lower computational cost showcase its potential and application prospects.
+    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news.
+Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption.
+Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
 2 Data-Copilot
-Data-Copilot is an LLM-based agent that accomplishes automated management, invocation, and processing of a large number of data by self-design sophisticated interface tools.
-Besides, Data-Copilot adeptly comprehends user intentions, and invokes well-designed tool interfaces for request solving, significantly reducing the need for tedious labor.
-As shown in  Figure 2, the LLM plays the role of a designer (Section 2.1) to independently design various tool interfaces to for data management and utilization.
-Besides, LLM also acts as a dispatcher (Section 2.2), invoking generated interfaces and deploying the workflow for problem-solving.
-Through the integration of two stages, Data-Copilot manages to efficiently handle a large volume of data-related tasks and caters to various user needs.
-2.1 Interface Design
-Just as Peter Naur pointed out, designing a set of versatile interface tools that are easy to manage and use is critical.
-Therefore, Data-Copilot designs a plethora of interfaces as tools, where each interface is composed of natural language (functional description) and code (implementation), responsible for data acquisition, processing, and visualization.
-As shown in Figure 2, beginning with some seed requests, LLM autonomously mimics and generates a large number of requests.
-Then LLM designs common interfaces for these requests, representing a set of universal needs.
-By gradually refining and expanding their functionalities, these interfaces progressively cover a more diverse array of needs.
-Ultimately, the LLM translates these interfaces into executable and callable code modules.
-Explore Data by Self-Request
-The design of the interface depends on two aspects: what kind of data is available and also what kind of demand the user proposes.
-Inspired by (Wang et al., 2022a; Dibia, 2023), Data-Copilot first autonomously explores the data to mine more requests.
-Then it proceeds to design the interfaces for these requests.
-Specifically, we generate a parsing file for each data source to help LLM understand the data.
+Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent.
+Data-Copilot operates in two phases: Data Exploration and Workflow Deployment.
+In the first phase (Section 2.1), Data-Copilot designs a self-request process to discover numerous potential needs based on data schemas (Step 1: exploration).
+Then based on these "self-instruct" requests, it generates code and abstracts universal interface modules for each request (Step 2: design).
+After that, it optimizes similar interfaces and reflects on the code compiler feedback, ensuring each interface is universal and correct (Step 3: optimization).
+Finally, Data-Copilot acquires many generic, error-free interfaces for subsequent use.
+The whole exploration-design-optimization process is operated in advance.
+When faced with real-time requests (Section 2.2), Data-Copilot adeptly comprehends user intentions, and invokes pre-designed interfaces for request solving, significantly reducing the need for tedious labor.
+We provide a detailed prompt for two phases in Section F.1 and F.2.
+2.1 Data Exploration
+Let’s review how human data analysts operate.
+Initially, they need to observe what data is available, understand the data formats, and determine how to use them.
+Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability.
+Similar to humans, Data-Copilot aims to explore data and derive insights from vast data sources.
+These insights include understanding the relations between the data, discovering the potential requests associated with the data, and organizing the code framework efficiently.
+Based on this, Data-Copilot summarizes these exploratory insights into many reusable code components and optimizes its performance.
+It significantly reduces the code complexity for real-time responses, thus improving the success rate.
+Request Exploration
+To explore data and mine insights, we design a self-request process.
+Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario.
+This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs must generate requests specifically related to these data.
+As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years.
+Each request involves one or more types of data.
+To achieve this, we first generate a parsing file for each data source to help LLM understand the data.
 Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example.
-This process does not require too much labor, as data providers usually offer a description of the data and access methods.
-Then we adopt a self-request process to explore data, i.e.
-the parsing files and a few seed requests are fed into the LLM and then LLM is instructed to generate more diverse requests by mimicking.
-This process aims to explore an extensive array of data sources while covering diverse user requests.
-In Figure 2, LLM generates four new requests based on the two seed requests and data parsing files.
-The detailed prompts are shown in Section F.1.
-Interface Definition
-In this step, Data-Copilot creates various interface tools to fulfill the previously generated requests.
-Specifically, we feed all data parsing files and all interfaces stored in the interface library (empty at first iteration) into LLM as a prompt.
-LLM needs to resolve each request one by one.
-When addressing each request, LLM can use existing interfaces from the interface library or redefine a new interface to fulfill the task.
-Concretely, if existing interfaces are insufficient to address the current request, LLM can create a new interface: Interface1={Name:"", Function description:"", Input and Output:""}...
-After that, LLM proposes a feasible solution using these self-design or existing interfaces: Solution="These three interface functions will meet your needs.
-You should first use the interface: {getBankData} to acquire data, then...".
-The prompts and cases are shown in Sections F.1 and G.1.
-Importantly, Data-Copilot only uses natural language to define the interface’s functions and their arguments, without considering the specific implementation, which is grammar-free.
-This process is similar to software architecture design by a human programmer.
-It allows Data-Copilot to focus on designing the layout and relation of the interfaces with different functionalities, rather than complying with programming grammar.
-Interface Merging
-To make the interface more universal, Data-Copilot considers whether each newly designed interface can be merged with the existing ones in the library.
-Specifically, when a new interface is designed, Data-Copilot also checks whether this interface is similar to the previous ones, in terms of functionality, arguments, etc.
-Two similar interfaces are merged to create a new and generalized interface.
-As shown in Figure 2, two interfaces {Interface: Query-GDP} and {Interface: Query-CPI} are merged into one {Interface: Query-GDP-CPI}.
-This merging process is similar to software developers wrapping similar modules into a new one.
-After this process, a large number of similar interfaces are merged. Each interface within the library possesses distinct functionalities, which are adeptly suited for real-time response applications.
-Interface Implementation
-Data-Copilot iteratively processes each request through interface definition and interface merging.
-Eventually, when all the requests can be satisfied by these self-design interfaces, Data-Copilot transforms these interface tools into executable and callable code modules.
-The whole interface design process is offline.
-As shown in Figure G3, Data-Copilot automatically produces five types of interfaces in the interface library: data acquisition, processing, prediction, visualization, and DataFrame manipulation.
-It transforms repetitive and tedious labor into an automated process, and also can effortlessly add new interfaces for additional requests or new data sources.
-Besides, Data-Copilot easily switches to other programming platforms and databases by simply re-generating the implementation code for the interface, demonstrating excellent scalability.
-2.2 Interface Dispatch
-In the first stage, Data-Copilot designs a variety of generic interface tools for data acquisition, processing, and visualization.
-These interfaces are versatile and easy to use, improving the efficiency of real-world response.
-These interfaces are invoked in this phase and form an efficient workflow for real-time requests (Figure 2).
-A detailed prompt and cases are shown in Sections F.2 and G.2
+Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to generate more diverse requests based on these data, covering different demand scenarios as comprehensively as possible.
+The brief prompt is shown as follows:
+###Instruction: Given some data and its description, please mimic these seed requests and generate more requests.
+The requests you generate should be as diverse as possible, covering more data types and common needs.
+###Seed Request: {request1, request2,...}
+###Parsing file for GDP_Data:{
+Description: This data records China’s annual and quarterly GDP...,
+Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
+Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
+Usage:{
+Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
+First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
+###Parsing file for Stock_data:{...},
+How to ensure the quality of the generated requests?
+A common issue is that the LLM often proposes a request about the data that doesn’t exist.
+To address this, we design a backward verification process to check the rationality of these requests.
+Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests.
+Upon manually check some requests, we find that the generated requests generally met our expectations.
+Interface Design
+After generating massive requests through the self-request, Data-Copilot considers abstracting universal interfaces from these requests.
+First, we need to define what an interface is in our paper.
+Similar to human programming practices, we refer to a function module abstracted from the code as an interface.
+Each interface is composed of a name, parameters, a functional description, and the implementation code, performing specific functions such as data retrieval, computing, and plotting.
+To abstract general interfaces from all requests, we design an interwoven process, performing interface design and interface optimization alternately for each request.
+Firstly, we feed related data parsing files and existing interfaces stored in the interface library (empty at first iteration) into LLMs as a prompt.
+LLMs need to generate a complete code for each request.
+When generating, LLMs are instructed to prioritize using the existing interfaces from the previous iterations.
+If these interfaces are insufficient to address the request, the LLMs design a new interface.
+As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...).
+As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
+Interface Optimization
+To improve the quality of interfaces, Data-Copilot optimizes them in two ways: merging similar interfaces to make them more universal and reflecting on compiler feedback to reduce errors.
+This optimization process is intertwined with the interface design for each request.
+Firstly, similar to software developers wrapping similar code modules to reduce redundancy, Data-Copilot considers whether each newly designed interface can be merged with the previous interfaces.
+LLMs are instructed to check the functionality, parameters, and logic of the new and existing interfaces, and merge those similar interface modules to create a new one.
+As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...).
+This interface optimization process makes each interface more general and rational.
+Besides, to reduce code hallucinations or formatting issues, Data-Copilot also self-reflects on the feedback of code execution for each request if compilation errors occur.
+We have LLMs perform self-reflection until the code passes or the maximum retry number is reached.
+This process ensures the correctness and usability of each interface.
+For each request, Data-Copilot alternates the above two steps: interface design and optimization until all the requests can be satisfied by these interfaces.
+Finally, as shown in Figure G3, Data-Copilot produces five types of interfaces: data acquisition, processing, prediction, visualization, and DataFrame manipulation.
+We provide some detailed cases in Section G.1.
+Besides, to handle newly added data sources, Data-Copilot can repeat the aforementioned exploration-design-optimization process on new data.
+It can create some new interface modules for these new data.
+2.2 Workflow Deployment
+In the first phase, Data-Copilot designs many generic interfaces with different functionalities.
+They are versatile and easy to use, simplifying the code generation of real-time responses.
+As shown in Figure 2, when faced with human requests, Data-Copilot only needs to invoke related interfaces and form an interface workflow.
+A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex call flow diagram.
+This workflow transforms raw data into visualized results.
+The detailed prompts and cases are shown in Sections F.2 and G.2.
 Intent Analysis
 To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks.
-For example, if the question is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2017 to May 2023 for comparison".
-To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this external information into LLM along with the original request to generate the parsed result.
-Planing Workflow
-Once the user’s intent is accurately understood, Data-Copilot plans a reasonable workflow to process the user’s request.
-We feed the user request and related interface descriptions into LLMs for the automatic deployment of workflows.
-Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot, based on the user’s request and the interface.
-To achieve this, Data-Copilot generates a response in JSON format representing each step of the scheduling: ###Interface Call: {step1={"arg":"", "function":"", "output":"" ,"description":""}, step2={..}, ..}.
-Please refer to Section F.2 for a detailed prompt.
-Data-Copilot meticulously orchestrates the scheduling of the interface in either a sequential or parallel manner.
-For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot first plans five steps in its workflow.
-In the first two steps, it dispatches {Data Acquisition}, {Calcuate Return} in parallel for three indices.
-The last three steps successively invoke the {Visualization} to sequentially plot three indices on the same canvas.
-In the second case, Data-Copilot deploys a loop workflow to sequentially calculate the indicators of 50 constituent stocks by calling the {Loop and Rank}.
+For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison".
+To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
+Planning Workflow
+After providing the user request and a description of the relevant interfaces, Data-Copilot plans a logical interface workflow to process the user request.
+Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot.
+The LLMs are instructed to output in JSON format for each step invocation: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
+For different requests, Data-Copilot designs interface workflows with different structures, such as sequential, parallel, or loop.
+For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot plans a five-step workflow.
+In the first two steps, it dispatches Data Acquisition(), Calcuate Return() in parallel for three indices.
+The last three steps successively invoke the Visualization() to sequentially plot three indices.
+In the second case, it deploys a loop workflow by calling the Loop().
 Multi-form Output
-Upon the deployment and execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text.
+Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text.
 Additionally, it also provides a comprehensive summary of the entire workflow.
-This systematic summary not only provides clear and concise results but also sheds light on the steps taken to achieve those results, thereby enhancing transparency and understanding of user requests.
+It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect.
 As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...".
-Data-Copilot first interprets the user’s intent based on local time.
-Then, it deploys a three-step workflow: 1) Invoking {get-GDP-data} interface to acquire historical GDP data.
-2) Invoking {predict-next-value} interface for forecasting.
+Data-Copilot first interprets the user’s intent.
+Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data.
+2) Invoking predict-next-value() interface for forecasting.
 3) Visualizing the output.
-3 Qualitative Evaluation
-3.1 Environment and Data Sources
-We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
-To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies.
-The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance.
-In the first phase, we use the gpt-4-0613222https://platform.openai.com/ for interface design and gpt-3.5-turbo-0613 for interface dispatch.
-3.2 Seeds and Datasets
-Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each.
-These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate.
-2) involves different levels of difficulty.
-Then we filter out highly similar requests by computing the sentence embedding of each request.
-Lastly, we retain 173 high-quality requests, which are used as the seeds for Interface Design phase of the Data-Copilot.
-Self-Request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request strategy to expand the request set.
-Specifically, we feed 173 manually proposed seed requests along with all data descriptions and schema into GPT-4.
-We instruct GPT-4 to generate massive requests, covering diverse data.
-In this phase, we generate a total of 6480 requests.
-Then we compute the semantic similarity between all requests and filter out those with high similarity.
-Lastly, we retain 3547 requests.
-We randomly select 3000 requests from the total set for subsequent interface design.
-Meanwhile, the remaining 547 requests are treated as a test set.
-We manually annotate the final answers, task types, and workflow types for each testing instance.
-The statistics of three dataset are shown in Tables 1 and C1.
-3.3 Qualitative Results
-The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request.
-The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces.
-For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
-Data-Copilot designs many versatile interface tools through a self-request.
-As shown in Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, dataframe manipulation, and visualization.
-These diverse interfaces endow the LLM with the capability to address complex requests.
-For instance, get-stock-prices-data interface integrates the functions of acquiring data in daily, weekly, and monthly frequencies, controlled by argument: freq, which is very similar to the process of programmer development.
-Data-Copilot deploys workflows with different structures.
-Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests.
-As shown in Figures G4 and G5, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface.
-It iteratively calculates the financial indicators for each stock.
-Please refer to Appendix E for more visualization cases.
-4 Quantitative Evaluation
-We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
 Table 1: Left: Statistics on human-proposed requests and self-request dataset. Right: Detail Statistics on different request types and workflow structures.
@@
 Total 237 184 126 547
+3 Qualitative Evaluation
+3.1 Environment and Data Sources
+We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data.
+To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies.
+The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance.
+In the first phase, we use the gpt-4-0613 for data exploration design and gpt-3.5-turbo-0613 for workflow deployment.
+3.2 Seeds and Datasets
+Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each.
+These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate.
+2) involves different levels of difficulty.
+Then we filter out highly similar requests by computing the sentence embedding of each request.
+Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
+en Self-request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request to expand the request set.
+Specifically, we feed 173 human-proposed requests along with all data descriptions and schema into GPT-4.
+We instruct GPT-4 to generate massive requests.
+In this phase, we generate 6480 requests, then compute the semantic similarity between all requests, and filter out those with high similarity, and retain 3547 requests.
+We randomly select 3000 requests from the total set for subsequent interface design.
+Meanwhile, the remaining 547 requests are treated as a test set.
+We manually annotate the final answers, task types, and workflow types for each testing instance.
+The statistics of three datasets are shown in Tables 1 and B1.
+3.3 Qualitative Results
+The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request.
+The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces.
+For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
+Data-Copilot designs versatile interfaces through data exploration.
+In Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, manipulation, and visualization.
+These diverse interfaces endow the LLM with the capability to address complex requests.
+For instance, get-stock-prices-data() interface integrates the functions of acquiring data by day, week, and month, controlled by argument: freq, which is similar to the process of programmer development.
+Data-Copilot deploys workflows with different structures.
+Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests.
+As shown in Figure 2, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface.
+It iteratively calculates the financial indicators for each stock.
+Please refer to Appendix E for more visualization cases.
+4 Quantitative Evaluation
+We quantitatively assess Data-Copilot in processing real-time user requests, where it could invoke the interfaces designed in the first phase.
 4.1 Experiments Settings
 Baselines 1.
-Global Planning (GP) (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests.
+Vanilla LLM (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests.
 2.
@@
 All methods are evaluated on the 547 testing instances (Section 3.2) three times.
-A detailed prompts of baselines is shown in Section F.3.
+The detailed prompts of all baselines are shown in Section F.3.
 Metrics We design three metrics for solution code and answer: Exec@1 Pass@1 and Efficiency.
@@
 Lastly, we calculate the total number of tokens consumed by the input prompt and output response, which denotes the solving efficiency of each method.
-Table 2: All methods use gpt-3.5 as based LLM.
+Table 2: We demonstrate the accuracy and efficiency of all methods with gpt-3.5.
 Methods Exec@1(%) Pass@1(%) #Token
-GP 50.8 28.5 823.6
-ReAct 63.4 44.1 1515.9
-Reflexion 75.7 59.0 2463.5
-Multi-Agent 73.6 57.4 2835.2
-Data-Copilot 79.9 70.2 561.2
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 3: All samples are divided into three subsets based on its complexity. We compute Pass@1.
 Methods Sequential Parallel Complex Overall
-GP 42.4% 29.0% 1.6% 28.5%
+Vanilla 42.4% 29.0% 1.6% 28.5%
 ReAct 56.3% 48.6% 14.6% 44.1%
@@
 Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%.
-In most cases, our system merely needs to invoke pre-designed interface tools and generate corresponding arguments to address user requests, significantly reducing the risk of errors.
-In contrast, baseline strategies must construct all code from scratch, which makes it easy to overlook certain details or make mistakes.
-Besides, the difference between Pass@1 and Exec@1 also indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
+In most cases, it merely needs to invoke pre-designed interface modules and generate corresponding arguments to address user requests, notably reducing the risk of errors.
+In contrast, baseline strategies must construct all code from scratch, making it easy to overlook certain details or make mistakes.
+Besides, the difference between Pass@1 and Exec@1 indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
 Data-Copilot reduces repetitive overhead for common requests.
 The consumed tokens represent computational overhead.
-Data-Copilot only consumes 561 tokens, which is just 30% of the consumption compared to baselines.
+Data-Copilot only consumes 561 tokens, just 30% of the baseline’s consumption.
 Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments.
-The primary token consumption lies in the input prompt which contains related tool descriptions.
-In contrast, baseline strategies may consume a substantial amount of tokens since they need to output a complete solving code for each request.
+The primary token consumption lies in the input prompt which contains related interfaces descriptions.
+In contrast, baseline strategies may consume a lot of tokens since they need to output a complete solving code for each request.
 In the real world, numerous requests are similar or even repetitive.
-Data-Copilot can reduce repetitive overhead by abstracting common demands into interface tools.
-However, baseline methods are compelled to repetitively generate code for each request.
+Data-Copilot can reduce repetitive overhead by abstracting common demands into interfaces.
+However, baseline methods have to repetitively generate code for each request.
 4.3 Analysis
@@
 To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex.
+Complex workflows mean some intricate cases, mainly involving loops or nested workflows.
 As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement.
@@
 We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing.
-In contrast, our methods effectively handle these tasks by invoking versatile interfaces and deploying workflow.
+In contrast, our methods can effectively handle these tasks by invoking versatile interfaces and deploying workflow.
 These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
@@
 We analyze the performance of Data-Copilot with different LLMs.
-As shown in Figure 4, we test GPT family APIs and various open-source models.
+In Figure 4, we test GPT family and various open-source models.
 We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results.
@@
 Specifically, we collect the solutions of Data-Copilot and Reflexion.
-Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix C).
+Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix B).
 We present results in Figure 4.
@@
 Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%.
-Besides, as for conciseness, Data-Copilot lags slightly behind its competitor, with 27.8% wins against 30% losses.
-This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are quite universal, accommodating a wide array of functions, which makes our system less concise.
+Besides, as for conciseness, Data-Copilot lags behind its competitor, with 27.8% wins against 30% losses.
+This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are universal, accommodating a wide array of functions, which makes our system less concise.
 The results of human evaluation are consistent with the previous assessment.
 Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
-Figure 3: Results of Data-Copilot with various LLMs.
+4.5 Performance of other programming languages
+Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces.
+We tested three programming languages: Python, C++, and Matlab.
+The results indicate that Python performed the best (Pass@1: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results.
+Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable.
+We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus.
+So ultimately, we opted for Python for Data-Copilot.
+Figure 3: Results of Data-Copilot using various LLMs as based model.
 LLM Pass@1 (%) #Token
@@
 5 Conclusion
-We propose a universal framework, Data-Copilot, to address a wide array of data-related tasks.
-It acts as a bridge between numerous heterogeneous data and humans, effectively managing, processing, and presenting data according to human tastes, reducing the dependence on tedious labor and expert knowledge.
-Acting like an experienced expert, Data-Copilot autonomously designs universal interface tools for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
-6 Acknowledgement
-This work is supported by the National Natural Science Foundation of China (No.
-62376245), the Key Research and Development Program of Zhejiang Province, China (No.
-2024C01034), the project of the Donghai Laboratory (Grant no.
-DH-2022ZY0013), and MOE Engineering Research Center of Digital Library.
+We propose a code-centric agent, Data-Copilot, to address a wide array of data-related tasks.
+Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests.
+It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
+Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
 Appendix
-Appendix A Discussion
-A.1 What is Data-Copilot?
-  • Data-Copilot is an efficient LLM-based data management agent, that processes massive data and rapidly responds to real-time human requests.
-  • It contains interface design and invocation phases. Interface design involves a self-exploration phase for data sources to create many versatile interfaces.
-  • For interface invocation, it parses human real-time requests and invokes the designed interfaces to retrieve and process relevant data and generates multi-form outputs, e.g., text, charts, and tables.
-A.2 Why Data-Copilot is designed in two stages?
-  • Reduces redundant generation: People often have a large number of similar or repetitive query needs, e.g., many people are interested in seeing which stock has the largest gain today.
-For these repetitive query requests, employing the LLM to write code from scratch for each user request would be highly inefficient and a significant waste of computational resources.
-To avoid inefficient and repetitive code generation for numerous similar requests, we abstract common demands into generic interfaces through data self-exploration in advance (Phase-1).
-For real-time requests (Phase-2), it only needs to invoke these interfaces, reducing redundant generation.
-  • Performs better and consumes fewer tokens: Compared to a complete code, it only needs to invoke the relevant interfaces, which is simpler and less prone to errors.
-Our results show that it performs better and consumes fewer tokens.
-  • Stronger interpretability: Compared to complex code, interface workflow is easier for humans to read and inspect.
-Appendix B Experiments Details
-During the interface design phase, we instruct LLM to generate Python code for each interface in the interface implementation.
-We filter out interfaces that cannot run properly at each iteration.
-For the interface Dispatch phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface.
+Appendix A Experiments Details
+During the data exploration phase, we instruct LLM to generate Python code for each interface.
+Then we abstract interfaces from these requests.
+To merge the similar interfaces for each iteration, we adopt the following strategy:
+1. Firstly, we calculate the semantic similarity between the names of all existing interfaces and the newly designed interface, then sort them by similarity from highest to lowest.
+2.
+Then, we feed the top-10 existing interfaces into the LLM, and instruct the LLM to sequentially check whether they are functionally similar to the new interface, or can be merged using new parameters.
+3. Lastly, we output the merged interface. If the LLM deems no merging necessary, the new interface is retained.
+Besides, for the workflow deployment phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface.
 We set the temperature to 0.2 for more consistent results.
-Appendix C Human Evaluation Details
+Appendix B Human Evaluation Details
 We collect the solutions of Data-Copilot and Reflexion on 173 human-proposed requests.
@@
 Experts are required to evaluate which of the two solutions is better based on these principles, and then we calculate the win rate of Data-Copilot against the Reflexion.
-Figure C1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
+Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
+Appendix C Limitations
+Data-Copilot proposes a new paradigm for addressing the data-related task, through LLM. But we want to highlight that it still remains some limitations or improvement spaces:
+1) Online Design Interface.
+The essence of Data-Copilot lies in effective interface design, a process that directly affects the effectiveness of subsequent interface deployments.
+Currently, this interface design process is conducted offline.
+Therefore, it is crucial to explore how to design the interface online and deploy it simultaneously.
+It greatly broadens the application scenarios of Data-Copilot.
+2) System stability The interface deployment process can occasionally be unstable.
+The main source of this instability is because LLM is not fully controllable.
+Despite their proficiency in generating the text, LLMs occasionally fail to follow the instructions or provide incorrect answers, thus causing anomalies in the interface workflow.
+Consequently, finding methods to minimize these uncertainties during the interface dispatch process should be a key consideration in the future.
+3) Data-Copilot possesses the potential to handle data from other domains effectively.
+Currently, our focus is on developing an LLM-based agent for the data domain.
+Due to the limited access to data, we chose the China Financial Data, which includes stocks, futures, finance, macroeconomics, and financial news.
+Although these data all belong to the financial domain, the data volume is extremely large, and the data schemas are highly different.
+The corresponding user’s requests are also diverse, which poses a great challenge to the current LLM.
+However, Data-Copilot has adeptly accomplished this task.
+Therefore, we believe Data-Copilot also possesses the potential to effectively handle data from other domains.
 Appendix D Related Works
-In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; OpenAI, 2023b; Wei et al., 2022b; Wu et al., 2023c) have revolutionized the field of natural language processing (NLP).
+In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa [Brown et al., 2020, Chowdhery et al., 2022, Zhang et al., 2022, Zeng et al., 2023, Touvron et al., 2023, Ouyang et al., 2022, OpenAI, 2023, Wei et al., 2022a] have revolutionized the field of natural language processing (NLP).
 These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning.
-The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism (Ouyang et al., 2022; Wang et al., 2022b; a).
+The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism [Ouyang et al., 2022, Wang et al., 2022b, a].
 LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs).
-Tool-enhanced studies  (Schick et al., 2023; Gao et al., 2022a; Qin et al., 2023a; Hao et al., 2023; Qin et al., 2023b) integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools.
-Several researchers have extended the scope of LLMs to include the other modality (Wu et al., 2023a; Surís et al., 2023; Shen et al., 2023; Liang et al., 2023b; Huang et al., 2023).
-In addition, there are many LLM-based agent applications (Xie et al., 2023), such as CAMEL (Li et al., 2023a), AutoGPT [^3] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^4] https://github.com/reworkd/AgentGPT, BabyAGI [^5] https://github.com/yoheinakajima/babyagi, BMTools [^6] https://github.com/OpenBMB/BMTools, LangChain [^7] https://github.com/hwchase17/langchain, Agentverse (Chen et al., 2023b), Autoagent (Chen et al., 2023a), MetaGPT (Hong et al., 2023), AutoGEN (Wu et al., 2023b), etc.
+Tool-enhanced studies  [Schick et al., 2023, Gao et al., 2022, Qin et al., 2023a, Hao et al., 2023, Qin et al., 2023b] integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools.
+Several researchers have extended the scope of LLMs to include the other modality [Wu et al., 2023a, Surís et al., 2023, Shen et al., 2023, Liang et al., 2023b, Huang et al., 2023].
+In addition, there are many LLM-based agent applications [Xie et al., 2023], such as CAMEL [Li et al., 2023a], AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse [Chen et al., 2023b], Autoagent [Chen et al., 2023a], MetaGPT [Hong et al., 2023], AutoGEN [Wu et al., 2023b], etc.
 Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks.
-Except for learning to operate the tools, several contemporaneous studies (Cai et al., 2023; Qian et al., 2023) have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning.
+Except for learning to operate the tools, several contemporaneous studies [Cai et al., 2023, Qian et al., 2023] have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning.
 These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
 Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers.
-FLAME (Joshi et al., 2023) investigates the feasibility of using NLP methods to manipulate Excel sheets.
-StructGPT (Jiang et al., 2023) explore reasoning abilities of LLM over structured data.
-LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration.
-Besides, many reseraches (Liu et al., 2023; Chang & Fosler-Lussier, 2023; Dong et al., 2023), like Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DAIL-SQL (Gao et al., 2023), DIN-SQL (Pourreza & Rafiei, 2023), DB-Copilot (Wang et al., 2023b), MAC-SQL (Wang et al., 2023a), ChatDB (Hu et al., 2023), SQL-PaLM (Sun et al., 2023), and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL and table rasoning.
-Chain-of-Table (Wang et al., 2024) proposes a step-by-step reasoning strategy based on the table.
+FLAME [Joshi et al., 2023] investigates the feasibility of using NLP methods to manipulate Excel sheets.
+StructGPT [Jiang et al., 2023] explore reasoning abilities of LLM over structured data.
+LiDA [Dibia, 2023] and GPT4-Analyst [Cheng et al., 2023] focus on automated data exploration.
+Besides, many reseraches [Liu et al., 2023, Chang and Fosler-Lussier, 2023, Dong et al., 2023], like Sheet-Copilot [Li et al., 2023b], BIRD [Li et al., 2023c], DAIL-SQL [Gao et al., 2023], DIN-SQL [Pourreza and Rafiei, 2023], PET-SQL [Li et al., 2024b], DB-Copilot [Wang et al., 2023b], MAC-SQL [Wang et al., 2023a], ACT-SQL [Zhang et al., 2023], ChatBI [Lian et al., 2024], CodeS [Li et al., 2024a], SQLPrompt [Sun et al., 2023a], ChatDB [Hu et al., 2023], SQL-PaLM [Sun et al., 2023b], and DB-GPT [Xue et al., 2023], EHRAgent [Shi et al., 2024] apply LLMs to Text2SQL and table rasoning.
+Chain-of-Table [Wang et al., 2024] proposes a step-by-step reasoning strategy based on the table.
 Distinct from these approaches, our Data-Copilot system is different in the following aspects:
@@
 Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
-(3) Distinct Implementation Approach: Data-Copilot designs its tools for data processing, rather than relying on SQL or pre-existing tools.
+(3) Distinct Implementation Approach: Data-Copilot is a code-centric agent for data processing, rather than relying on SQL or pre-existing tools.
 This approach is more flexible and can meet complex user requirements.
@@
 In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas.
-These concise workflows can cope with such complex requests well, which suggests that the interface design and dispatch process of Data-Copilot are rational and effective.
+These concise workflows can cope with such complex requests well, which suggests that the data exploration and workflow deployment process of Data-Copilot are rational and effective.
 Diverse Sources
@@
 We provide detailed prompts for our Data-Copilot and baselines.
-Specifically, for the interface design, we provide detailed prompts in Section F.1 for four phases: Self-Request, Interface Definition, Interface Merging, and Interface Implementation.
-For the interface dispatch, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow.
-Additionally, we outline prompts for all baselines in Section F.3: Global Planning, ReAct, Reflection, and Multi-Agent Collaboration strategies.
-F.1 Prompts For Interface Design
+Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-request, interface design, and interface optimization.
+For the workflow deployment phase, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow.
+Additionally, we outline prompts for all baselines in Section F.3: Vanilla LLM, ReAct, Reflection, and Multi-Agent Collaboration strategies.
+F.1 Prompts For Data Exploration
 Explore Data by Self-Request Phase
@@
 ###Seed Request: {request1, request2,...}
-###Data files:{
-GDP-Data:{
+###parsing file for GDP-Data:{
 Description: This data records China’s annual and quarterly GDP...,
@@
 Last Rows: {2018Q4 900309.5 6.60, ...}}}
-stock_basic:...,
-Interface Definition Phase:
-###Instruction: You are an experienced program coder.
-Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces (similar to pseudocode) to resolve my request.
+###parsing file for stock_data: {...}
+Interface Design:
+###Instruction: You are an experienced program coder. Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces to resolve my request.
 (1) You should define the name, function, inputs, and outputs of the interface.
-Please describe the functionality of the interface as accurately as possible and do not write any implementation code in the new interface, just design the interface as existing interfaces.
+Please describe the functionality of the interface as accurately as possible and write complete implementation code in the new interface.
 (2) Finally please explain how to resolve my request using your newly designed interfaces or existing interfaces in the neural language.
@@
 ###The user request: {input request}
-###Data Files: {All data files}
+###Data Files: {All data parsing files}
 ###The existing interfaces: {all interfaces in library}
-Interface Merging Phase
+Interface Optimization
 ###Instruction: Please check that the interface you have designed can be merged with any existing interfaces in the library.
 (1) You should merge interfaces with similar functionality and similar input and output formats into a new interface.
-(2) You can use parameters to control the different inputs. Finally
+(2) You can use parameters to control the different inputs, if you want to merge two interfaces.
 (3) Please explain your reason for merging and output all interfaces in the library after merging.
@@
 The reasons for merging: {reason}
-All interfaces after merging: {interface1, interface2, merged interface3, ...}
-Interface Implementation Phase:
-###Instruction: Please generate the complete code according to the definition of a given interface and data access methods.
-The generated code must be in strict accordance with the definition of the interface and the formatting requirements of the arguments.
-###Interface: {Input Interface}
-###Data Files: {All data files}
-F.2 Prompts For Interface Dispatch
+Interfaces after merging: {interface1, interface2, optimized interface3, ..}
+F.2 Prompts For Workflow Deployment
 Intent Analysis Phase
@@
 F.3 Prompts For Baselines
-Global Planning
+Vanilla LLM
 ###Instruction: You are an artificial intelligence assistant.
@@
 Appendix G Case Study
-G.1 Cases for Interface design
----- The First Request: Interface Definition Phase -----
+Detailed cases of data exploration and workflow deployment are presented in Sections G.1 and G.2.
+G.1 Cases for Data Exploration
+---- The First Request: Interface Design -----
 Input Request: Get the daily opening price data of Bank of Hangzhou from Jan 23, 2018 to Mar 13, 2019 and plot the trend.
@@
 New Design Interfaces:
-{Interface1=Interface Name:getBankData, Function:.., Input:.., Output:..}
-{Interface2=Interface Name:prepareDataForPlotting, Function:.., Input:.., Output:..}
-{Interface3=Interface Name:plotData, Function:.., Input:.., Output:..}
+{Interface1=Name:getBankData(), Function:.., Input:.., Output:..}
+{Interface2=Name:prepareDataForPlotting(), Function:.., Input:.., Output:..}
+{Interface3=Name:plotData(), Function:.., Input:.., Output:..}
 The solving process for request: These three new interface functions will meet your needs.
 In your program, you would first get the data using getBankData, then process the data using prepareDataForPlotting, and finally use plotData to plot the trend.
+Complete code: {Complete code using three new interfaces...}
 ---- The Second Request: Interface Definition Phase-------
@@
 New Designed Interfaces:
-{Interface4=Interface Name: getStockData..}
+{Interface4=Name: getStockData()..}
 Existing Interfaces
-{Interface1=Interface Name:getBankData, ...}
-{Interface2=Interface Name:prepareDataForPlotting, ...}
-{Interface3=Interface Name:plotData, ...}
+{Interface1=Name:getBankData(), ...}
+{Interface2=Name:prepareDataForPlotting(), ...}
+{Interface3=Name:plotData(), ...}
 The solving process for request: You need to call these functions to meet your needs.
-First, use new interface {getStockData} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting} to process this data, and finally use {plotData} to plot the trend...
----- Interface Merge Phase -------
+First, use new interface {getStockData()} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting()} to process this data, and finally use {plotData()} to plot the trend...
+Complete code: {Complete code using three new interfaces...}
+---- Interface Optimization Phase -------
 The reason for merging: Yes, these interface functions can be merged and reformulated to eliminate duplicate functions.
-We can merge getBankData and getStockData into a more generic function get-stock-data, which can fetch any financial data, including bank and stock data.
+We can merge getBankData() and getStockData() into a more generic function get-financial-data(), which can fetch any financial data, including bank and stock data.
 In addition, we can keep the prepareDataForPlotting and plotData functions, as they can also be applied to different types of financial data]
-{Interface1=Interface Name:get-stock-data, ...}
-{Interface2=Interface Name:prepareDataForPlotting, ...}
-{Interface3=Interface Name:plotData, ...}
----- Interface Implementation -------
-Using get_stock_data as examples:
-def get_stock_data(name:str=’’,start_d:str=’’,end_d:str=’’,freq:str)->pd.DataFrame:
-Retrieves the daily/weekly/monthly price data for a given stock code during a specific time period.
-Args:
--name (str): The name of the stock.
--start_d (str): The start date in the format YYYYMMDD’.
--end_d (str): The end date in YYYYMMDD’.
--freq (str): The frequency of the price, can be daily’, weekly’, or monthly’.
-Returns:
--pd.DataFrame: A dataframe that contains the daily/weekly/monthly data. The output columns contain stock_code, trade_date, open, high, low, close, pre_close, change), pct_chg,vol,amount
-stock_code = get_stock_code(name)
-if freq == daily’:
+{Interface1=Interface Name:get-financial-data(), ...}
+{Interface2=Interface Name:prepareDataForPlotting(), ...}
+{Interface3=Interface Name:plotData(), ...}
 G.2 Cases for Real-time Request
@@
 Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The contribution of the article is clearly stated in Abstract and Section 1 (Introduction), and is fully consistent with the experiment (Section 3).
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss limitations in Appendix Section C.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: This paper does not deal with the theoretical derivation part.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We have provided a code implementation for our system. The codebase has been open-sourced.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We have provided all the experimental details in the article (Section 4.1), prompt design (Appendix A, B, F.1 F.2 F.3).
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We provide details of the experiment, reporting the mean and standard deviation after multiple experiments in Table 2.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: In our experiments, we report the model of the API interfaces used and the tokens consumed for each experiment.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [N/A]
+  44. Justification: This paper does not involve any issues related to the Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: We propose an autonomous data analysis agent that does not generate any negative social impact.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [No]
+  54. Justification: Our method does not pose such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We utilized financial data provided by Tushare, which has been duly cited in the paper.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [No]
+  64. Justification: The paper does not release new assets
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [Yes]
+  69. Justification: We invited experts to annotate and evaluate the data, providing them with detailed instructions and guidelines.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our paper does not present these risks.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.